# Q-Frame: Query-aware Frame Selection and Multi-Resolution Adaptation for Video-LLMs

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant success in visual understanding tasks. However, challenges persist in adapting these models for video comprehension due to the large volume of data and temporal complexity. Existing Video-LLMs using uniform frame sampling often struggle to capture the query-related crucial spatiotemporal clues of videos effectively. In this paper, we introduce Q-Frame, a novel approach for adaptive frame selection and multi-resolution scaling tailored to the video’s content and the specific query. Q-Frame employs a training-free, plug-and-play strategy generated by a text-image matching network like CLIP, utilizing the Gumbel-Max trick for efficient frame selection. Q-Frame allows Video-LLMs to process more frames without exceeding computational limits, thereby preserving critical temporal and spatial information. We demonstrate Q-Frame’s effectiveness through extensive experiments on benchmark datasets, including MLVU, LongVideoBench, and Video-MME, illustrating its superiority over existing methods and its applicability across various video understanding tasks.
@@ 1 Introduction
Video understanding, however, presents unique challenges due to the inherently large amount of video frames and the continuous nature of video content. Despite progress in video understanding, current methods suffer from inherent limitations due to the temporal redundancy in videos and the restricted context length within MLLMs. A typical video contains a large number of frames—e.g., a 3-minute video at 24 frames per second (fps) results in approximately 4320 frames. Modern MLLMs remain constrained by their context—VideoLLaMA2 [5] supports only 2K tokens, while VILA-V1.5 [15] extends this to roughly 4K tokens. It becomes clear that using all frames for understanding is computationally infeasible. Previous approaches have relied on uniform frame sampling, where a fixed subset of frames is selected at regular intervals. While this method is straightforward, it has several drawbacks: it produces sparse frame selections that break temporal continuity, it follows a static and query-agnostic sampling strategy, and it treats all frames equally, which is inefficient and suboptimal for fine-grained tasks [14, 20, 39, 28].
@@
To address these limitations, we propose Q-Frame, a novel framework for query-adaptive frame selection and resolution adaptation in Video-LLMs. As illustrated in Figure 1, Q-Frame dynamically selects the most relevant frames from a video and adjusts their resolutions according to both the video content and the given query. In contrast to prior approaches that treat all frames uniformly, Q-Frame introduces a dynamic strategy that makes frame sampling query-aware and computationally efficient. This strategy is derived from a discrete probability distribution parameterized by the output of a text–image matching model (e.g., CLIP [22]). By leveraging the Gumbel-Max trick [8], we can sample from this distribution to obtain an optimal combination of frames without resorting to complex optimization or any additional training, making Q-Frame easy to implement. Importantly, as shown in Figure 2, Q-Frame allows a Video-LLM to process a greater number of informative frames without exceeding its context limit, thus preserving crucial visual details that would be lost with uniform sampling. Moreover, Q-Frame is model-agnostic and plug-and-play: It requires no fine-tuning of the Video-LLM and can be applied to both open-source models [27, 10] and closed-source API-based models [21] alike.
@@ 2 Related Work
Video Large Language Models. Early video understanding approaches primarily relied on handcrafted features or conventional deep networks with limited capacity to capture temporal dynamics and semantic richness in visual content [26]. Recent advances in Multimodal Large Language Models (MLLMs) have revolutionized video understanding. The Video-LLMs can be categorized into the following three categories. Firstly, VideoChat [35], LLaVA-Hound [38], InternVideo2 [28], ShareGPT4Video [3], Video-LLaVA [14], and LLaVA-Video [39] leverage synthesized video-instruction data (e.g., video captions and open-ended QA) to enhance fine-grained action understanding and narrative tracking. Secondly, Video-ChatGPT [20], MovieChat [25], TimeChat [23], Video-LLaMA [35], and Video-CCAM [6] develop efficient feature compression methods for holistic video semantics, employing temporal aggregation and contrastive spatial-temporal modeling. Finally, LongVA [37], LongVILA [32], and Kangaroo [19] extend the language model context windows to accommodate extended visual token sequences for long-form video analysis.
Keyframe Selection. Keyframe selection serves as a fundamental technique for identifying representative frame combinations that balance content coverage and redundancy reduction. In the context of early action recognition, Adaframe [31] pioneered reinforcement learning (RL)-based adaptive frame selection, later extended by [30] through multi-agent decision-making. Dynamic video analysis motivated motion-guided approaches like MGSampler [40] and its probabilistic variant PMGSampler [2], which prioritize motion-salient frames. Most Video-LLMs [14, 20, 39, 28] adopt query-agnostic uniform sampling, often resulting in incomplete relevant content capture. Recent advances introduce query-conditioned strategies. KeyVideoLLM [13] implements query-aware top-K semantic retrieval, whereas Frame-Voyager [33] employs loss-driven combinatorial optimization. Nevertheless, existing methods face persistent challenges in temporal fragmentation and rigid resolution constraints during frame representation.
Any-Resolution Processing. MLLMs typically encode visual inputs as discrete vision tokens for cross-modal alignment. Seminal works [16, 17] establish this paradigm through fixed-resolution vision encoders, inherently constraining input flexibility. Subsequent approaches [12, 4, 18] address native resolution preservation by partitioning images into fixed-grid patches yet retain structural rigidity and disregard spatial dependencies during tokenization. A breakthrough emerges with Qwen2-VL’s [27] Native Dynamic Resolution framework, enabling adaptive visual token generation across varying input resolutions. Although any-resolution techniques have been well-researched for Image-LLMs, the dynamic resolution adaptation of Video-LLMs is still blank.
@@ 3.1 Overview
Figure 3: The overall framework of Q-Frame. Q-Frame is composed of Cross-modal Query Retrival (CQR), Query-Aware Frame Selection (QFS), and Multi-Resolution Adaptation (MRA). CQR focuses on retrieving the most semantically relevant frames from the video based on the textual query, ensuring that only meaningful visual information is considered. QFS is designed to adaptively select frames based on their relevance to the query, enhancing efficiency by concentrating on the most important temporal segments. MRA aims to optimize computational resources by assigning varying resolutions to frames, preserving fine details in important frames while reducing costs for less critical ones. It should be noted that the preprocessing strategy of Video-LLMs is different, and the MRA in the dotted line is not applicable to every model.
@@ 3.2 Cross-modal Query Retrieval
where $\text{VLM}_{\text{vision}}(\cdot)$ and $\text{VLM}_{\text{text}}(\cdot)$ denote the vision and text encoders of the VLM, respectively. To compute the semantic relevance of each frame, we estimate similarity by the inner product of embeddings:
@@ 4.1 Experimental Setup
Baseline Models: To validate Q-Frame, as detailed in Table 1, we select representative open-source models, namely VILA-V1.5 [15] and Qwen2-VL [1], as well as the closed-source API-based GPT-4o [21]. Q-Frame is integrated with these baselines to assess its performance relative to recent state-of-the-art Video-LLMs. In order to maintain comparability with FRAME-VOYAGE [33], experiments using VILA-V1.5 employ QFS with fixed input frames and resolutions. Additionally, because of the closed-source nature of GPT-4o, the effect of its preprocessing on image resolution remains unknown. Consequently, the GPT-4o’s experiments only utilize QFS. Consequently, our evaluation includes two experimental settings: one with a fixed number of input frames and another with a fixed token budget.
@@ 4.3 Ablation Study
Multi-resolution Strategy of MRA: Table 5 highlights the effects of various resolution strategies within the MRA module. The results emphasize the advantages of multi-resolution processing, where allocating a high resolution to key frames and lower resolutions to less important frames enhances performance. The highest accuracy of 58.4% is achieved when all three resolution categories (high, medium, and low) are utilized, illustrating that a flexible resolution strategy is ideal. When only high-resolution frames are employed, the accuracy declines to 57.6%, and relying solely on medium or low resolutions leads to further performance degradation (55.9% and 49.0%, respectively). These results suggest that a balanced multi-resolution approach offers the best compromise between performance and computational efficiency.
@@ 4.4 Qualitative Analysis
Figure 5 illustrates a representative example from our qualitative analysis. In this example, the query "How do the players celebrate scoring the first goal in this video?" is answered correctly as "Kneeling on the field and screaming" when Q-Frame is applied, whereas uniform sampling fails to capture sufficient context. This comparison highlights that Q-Frame effectively selects key frames and adapts their resolutions to preserve critical temporal and spatial details, ultimately enabling more accurate and coherent responses. Additionally, with the same consumption level, our Q-Frame can process more frames and assign appropriate weights based on their significance.
@@ B.4 Additional Ablation Study
Overhead: Although Q-Frame doesn’t require additional training, it introduces preprocessing overhead in the inference phase. The ablation study of the overhead of Q-Frame is shown in Table 14. The overhead of Q-Frame focuses on calculating embeddings and performing sampling (including resolution allocation). As the embeddings of candidate frames can be computed in parallel, the computational cost of this part is relatively manageable. Additionally, as the number of candidate frames increases, the burden brought by MRA is almost negligible. Compared to Video-LLM’s processing time, Q-Frame can introduce about 5 times more effective frames with minimal impact on time.
Candidate Sapmled Latency(ms) LongVideoBench
@@ B.5 Additional Case Analysis
We present a comparison visualization of uniform sampling and our Q-Frame. As illustrated in Figures 6, 7, and 8, Q-Frame more accurately identifies the video frame where the answer is located, providing a more reliable source of information for Video-LLMs. Furthermore, due to the introduction of the Gumble-Max trick, the sampling in QFS continues over time, which solves the sparsity problem caused by sampling to some extent.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant success in visual understanding tasks. However, challenges persist in adapting these models for video comprehension due to the large volume of data and temporal complexity. Existing Video-LLMs using uniform frame sampling often struggle to capture the query-related crucial spatiotemporal clues of videos effectively. In this paper, we introduce Q-Frame, a novel approach for adaptive frame selection and multi-resolution scaling tailored to the video’s content and the specific query. Q-Frame employs a training-free, plug-and-play strategy based on a text-image matching network like CLIP, utilizing the Gumbel-Max trick for efficient frame selection. Q-Frame allows Video-LLMs to process more frames without exceeding computational limits, thereby preserving critical temporal and spatial information. We demonstrate Q-Frame’s effectiveness through extensive experiments on benchmark datasets, including MLVU, LongVideoBench, and Video-MME, illustrating its superiority over existing methods and its applicability across various video understanding tasks.
@@ 1 Introduction
Video understanding, however, presents unique challenges due to the inherently large number of video frames and the continuous nature of video content. Despite progress in video understanding, current methods suffer from inherent limitations due to the temporal redundancy in videos and the restricted context length within MLLMs. A typical video contains a large number of frames—e.g., a 3-minute video at 24 frames per second (fps) results in approximately 4,320 frames. Modern MLLMs remain constrained by their context—VideoLLaMA2 [5] supports only 2,000 tokens, while VILA-V1.5 [15] extends this to roughly 4,000 tokens. It becomes clear that using all frames for understanding is computationally infeasible. Previous approaches have relied on uniform frame sampling, where a fixed subset of frames is selected at regular intervals. While this method is straightforward, it has several drawbacks: it results in sparse frame selections that break temporal continuity, it follows a static and query-agnostic sampling strategy, and it treats all frames equally, which is inefficient and suboptimal for fine-grained tasks [14, 20, 39, 28].
@@
To address these limitations, we propose Q-Frame, a novel framework for query-adaptive frame selection and resolution adaptation in Video-LLMs. As illustrated in Figure 1, Q-Frame dynamically selects the most relevant frames from a video and adjusts their resolutions according to both the video content and the given query. In contrast to prior approaches that treat all frames uniformly, Q-Frame introduces a dynamic strategy that makes frame sampling query-aware and computationally efficient. This strategy is derived from a discrete probability distribution parameterized by the output of a text–image matching model (e.g., CLIP [22]). By leveraging the Gumbel-Max trick [8], we can sample from this distribution to obtain an optimal combination of frames without resorting to complex optimization or any additional training, making Q-Frame easy to implement. Importantly, as shown in Figure 2, Q-Frame allows a Video-LLM to process a greater number of informative frames without exceeding the computational budget, thus preserving crucial visual details that would be lost with uniform sampling. Moreover, Q-Frame is model-agnostic and plug-and-play, requiring no fine-tuning of the Video-LLM and can be applied to both open-source models [27, 10] and closed-source API-based models [21].
@@ 2 Related Work
Video Large Language Models. Early video understanding approaches primarily relied on handcrafted features or conventional deep networks with limited capacity to capture temporal dynamics and semantic richness in visual content [26]. Recent advances in Multimodal Large Language Models (MLLMs) have revolutionized video understanding. The Video-LLMs can be categorized into the following three categories. Firstly, VideoChat [35], LLaVA-Hound [38], InternVideo2 [28], ShareGPT4Video [3], Video-LLaVA [14], and LLaVA-Video [39] leverage synthesized video-instruction data (e.g., video captions and open-ended QA) to enhance fine-grained action understanding and narrative tracking. Secondly, Video-ChatGPT [20], MovieChat [25], TimeChat [23], Video-LLaMA [35], and Video-CCAM [6] develop efficient feature compression methods for holistic video semantics, employing temporal aggregation and contrastive spatial-temporal modeling. Finally, LongVA [37], LongVILA [32], and Kangaroo [19] extend the LLM’s context length to accommodate extended visual token sequences for long video analysis.
Keyframe Selection. Keyframe selection is a fundamental technique for identifying representative frame combinations that balance content coverage and reduce redundancy. In the context of early action recognition, AdaFrame [31] pioneered reinforcement learning (RL)-based adaptive frame selection, later extended by [30] through multi-agent decision-making. Dynamic video analysis has led to the development of motion-guided approaches such as MGSampler [40] and its probabilistic variant PMGSampler [2], which prioritize motion-salient frames. Most Video-LLMs [14, 20, 39, 28] adopt query-agnostic uniform sampling, often resulting in incomplete relevant content capture. Recent advances introduce query-conditioned strategies. KeyVideoLLM [13] implements query-aware top-K semantic retrieval, whereas Frame-Voyager [33] employs loss-driven combinatorial optimization. Nevertheless, existing methods face persistent challenges in temporal fragmentation and rigid resolution constraints during frame representation.
Any-Resolution Processing. MLLMs typically encode visual inputs as discrete visual tokens for cross-modal alignment. Seminal works [16, 17] establish this paradigm through fixed-resolution vision encoders, inherently constraining input flexibility. Subsequent approaches [12, 4, 18] address native resolution preservation by partitioning images into fixed-grid patches, yet retain structural rigidity and disregard spatial dependencies during tokenization. A breakthrough emerges with Qwen2-VL’s [27] Native Dynamic Resolution framework, enabling adaptive visual token generation across varying input resolutions. Although any-resolution techniques have been well-researched for Image-LLMs, the dynamic resolution adaptation of Video-LLMs remains unexplored.
@@ 3.1 Overview
Figure 3: The overall framework of Q-Frame. Q-Frame is composed of Cross-modal Query Retrieval (CQR), Query-Aware Frame Selection (QFS), and Multi-Resolution Adaptation (MRA). CQR focuses on retrieving the most semantically relevant frames from the video based on the textual query, ensuring that only meaningful visual information is considered. QFS is designed to adaptively select frames based on their relevance to the query, enhancing efficiency by concentrating on the most important temporal segments. MRA aims to optimize computational resources by assigning varying resolutions to frames, preserving fine details in important frames while reducing costs for less critical ones. It should be noted that the preprocessing strategy of Video-LLMs is different, and the MRA in the dotted line is not applicable to every model.
@@ 3.2 Cross-modal Query Retrieval
where $\text{VLM}_{\text{vision}}(\cdot)$ and $\text{VLM}_{\text{text}}(\cdot)$ denote the vision and text encoders of the VLM, respectively. To compute the semantic relevance of each frame, we compute the similarity as the inner product between the query and frame embeddings:
@@ 4.1 Experimental Setup
Baseline Models: To validate Q-Frame, as detailed in Table 1, we select representative open-source models, namely VILA-V1.5 [15] and Qwen2-VL [1], as well as the closed-source API-based GPT-4o [21]. Q-Frame is integrated with these baselines to assess its performance relative to recent state-of-the-art Video-LLMs. In order to maintain comparability with FRAME-VOYAGE [33], experiments using VILA-V1.5 employ QFS with fixed input frames and resolutions. Additionally, because of the closed-source nature of GPT-4o, the effect of its preprocessing on image resolution remains unknown. Consequently, the GPT-4o’s experiments only utilize QFS. Our evaluation includes two experimental settings: one with a fixed number of input frames and another with a fixed token budget.
@@ 4.3 Ablation Study
Multi-resolution Strategy: Table 5 highlights the effects of various resolution strategies within the MRA module. The results emphasize the advantages of multi-resolution processing, where allocating a high resolution to key frames and lower resolutions to less important frames enhances performance. The highest accuracy of 58.4% is achieved when all three resolution categories (high, medium, and low) are utilized, illustrating that a flexible resolution strategy is ideal. When only high-resolution frames are employed, the accuracy declines to 57.6%, and relying solely on medium or low resolutions leads to further performance degradation (55.9% and 49.0%, respectively). These results suggest that a balanced multi-resolution approach offers the best compromise between performance and computational efficiency.
@@ 4.4 Qualitative Analysis
Figure 5 illustrates a representative example from our qualitative analysis. In this example, the query "How do the players celebrate scoring the first goal in this video?" is answered correctly as "Kneeling on the field and screaming" when Q-Frame is applied, whereas uniform sampling fails to capture sufficient context. This comparison highlights that Q-Frame effectively selects key frames and adapts their resolutions to preserve critical temporal and spatial details, ultimately enabling more accurate and coherent responses. Additionally, with the same consumption level, our Q-Frame can process more frames and assign appropriate resolutions based on relevance.
@@ B.4 Additional Ablation Study
Overhead: Although Q-Frame doesn’t require additional training, it introduces preprocessing overhead in the inference phase. The ablation study of the overhead of Q-Frame is shown in Table 14. The overhead of Q-Frame focuses on calculating embeddings and performing sampling (including resolution allocation). As candidate frame embeddings can be computed in parallel, the computational cost of this part is relatively manageable. Additionally, as the number of candidate frames increases, the burden brought by MRA is almost negligible. Compared to Video-LLM’s processing time, Q-Frame can introduce about 5 times more effective frames with minimal impact on time.
Candidate Sampled Latency(ms) LongVideoBench
@@ B.5 Additional Case Analysis
We present a comparison visualization of uniform sampling and our Q-Frame. As illustrated in Figures 6, 7, and 8, Q-Frame more accurately identifies the video frame where the answer is located, providing a more reliable source of information for Video-LLMs. Furthermore, due to the introduction of the Gumbel-Max trick, the sampling in QFS continues over time, which solves the sparsity problem caused by sampling to some extent.